Consider grains of sand 100 µm in diameter. Calculate the number of grains and number of moles if one were to line the grains in a chain from earth to the moon, a distance of 384,000 km.
Answer #1

D = 100 microm = 100*10^-6 m

# grains to moon

D = 384km = 384*10^3 m

then

# = ( 384*10^3)/(100*10^-6) = 3840000000 grains

number of moles

1 mol = 6.022*10^23 grains

3840000000/(6.022*10^23) = 6.37661*10^-15 mol of grains